The weight of students ina college has a mean of 40kgs and standard dev of 5 kg.In a random sample of 5 students what is the probability that
1) all have weight>50kg
2) 3 will have weight>50kg
1
Expert's answer
2020-05-25T20:57:50-0400

Let "X" be the weight of a student, and random variable "Z" has standard normal distribution "Z" ~"N(0,1)" .

We assume that "X" has normal distribution, then "X" ~"N(40,5^2)", and

"\frac{X-40}{5}" ~"Z".

The probability that student weight ">50" is

"p=P(X>50)=P(\frac{X-40}{5}>\frac{50-40}{5})=\\
P(Z>2)=1-P(Z\leq2)=\\1-0.9772=0.0228."

Let "k" be the number of students weighing ">50", then "k" has a binomial distribution

with the probability of success "p=0.0228" and parameter "n=5".

a)

"p(k=5)=C_n^kp^k(1-p)^{n-k}=\\
C_5^5*0.0228^5*0.9772^0=0.0228^5\approx 6*10^{-9}."

b)

"p(k=3)=C_n^kp^k(1-p)^{n-k}=\\
C_5^3*0.0228^3*0.9772^2=\\
\frac{5*4*3}{1*2*3}*0.0228^3*0.9772^2\approx 10^{-4}."

Answer: a) "6*10^{-9}" , b) "10^{-4}."